# Stone Skipping Black Holes in Ultralight Dark Matter Solitons

### 02 Apr 2026 v2

@@ I Introduction
We find that stone skipping is driven by the dipole perturbations of the soliton. The dipole response of self-gravitating systems has long been recognized as a source of weakly damped oscillations, which are often referred to as seiche modes, in classical stellar dynamics [69, 70, 71, 63]. Our work extends this physical picture to the wave-mechanical context of ULDM via the eigenmode decomposition framework developed by Zagorac et al. [74]. Through controlled simulations with selectively excited or suppressed modes, we establish that dipole terms are both necessary and sufficient to trigger stone skipping. We then model this phenomenon as a resonance in a forced, damped oscillator. The time-dependent dipole provides periodic driving, while dynamical friction acts as dissipation. When the forcing frequency approaches the natural epicyclic frequency the system enters a resonance window where energy transfer from the soliton can overcome secular damping.
@@ III Eigenmode Decomposition
Our approach is based on Ref. [74] which we generalize to include the full set of angular states. The macroscopic ULDM wave function is equivalent to an effective single-particle wave function up to its normalization so we may construct eigenmodes using standard quantum-mechanical methods. In operator form the Schrödinger sector of the system is
@@ VII Conclusion
Several directions follow naturally. A first priority is to embed the soliton in a more realistic galactic environment by incorporating an outer halo and assessing the impact of ambient ULDM granule fluctuations and relaxation processes [9, 35] on any dipole excitation of the soliton. Second, the resonance model should be generalized to eccentric or inclined orbits and multi-body configurations. Third, given that unequal-mass binaries or mergers could excite a dipole term it will be valuable to delineate the parameter space in which soliton-driven resonance works against a drag-driven inspiral. Fourth, one could generalize this work to models with a mixture of cold and ultralight dark matter [57] and multi-component ULDM models [61, 33, 5, 31]. Finally, we did not describe the original excitation of the dipole (and other) modes by the orbiting black hole(s) within perturbation theory – a full treatment of this would be both interesting and useful.
@@ B.2 Linearised quasi-circular expansion
$\ddot{\eta}=\frac{1}{r_{0}^{2}}\,\Bigl[F_{\phi}(t)-2r_{0}\,\Omega_{0}\,\dot{\xi}-\gamma_{\phi}\,r_{0}^{2}\,\bigl(\Omega_{0}+\dot{\eta}\bigr)-2\gamma_{\phi}\Omega_{0}r_{0}\xi\Bigr].$ (114)
@@ B.3 Forced damped harmonic oscillator
$\Delta=\frac{16\bigl(2\Gamma^{2}-\kappa^{2}\bigr)^{2}}{9}-\frac{4\kappa^{4}}{3}=\frac{4\bigl(\kappa^{2}-8\Gamma^{2}\bigr)^{2}-48\,\Gamma^{4}}{9}\geq 0,$ (139)
@@
$\mu_{0}=0,\qquad\mu_{\pm}=\frac{2\kappa^{2}-4\Gamma^{2}\pm\sqrt{\kappa^{4}-16\kappa^{2}\Gamma^{2}+16\Gamma^{4}}}{3}.$ (141)

### 13 Apr 2026 v3

@@ I Introduction
We find that stone skipping is driven by the dipole perturbations of the soliton. The dipole response of self-gravitating systems has long been recognized as a source of weakly damped oscillations, which are often referred to as seiche modes, in classical stellar dynamics [70, 71, 72, 64]. Our work extends this physical picture to the wave-mechanical context of ULDM. More broadly, oscillations and random-walk behaviour of FDM soliton cores have been interpreted in terms of interference between the ground state and excited states in a fixed potential [42], while the specific perturbative eigenmode framework used here builds on Zagorac et al. [77]. Through controlled simulations with selectively excited or suppressed modes, we establish that dipole terms are both necessary and sufficient to trigger stone skipping. We then model this phenomenon as a resonance in a forced, damped oscillator. The time-dependent dipole provides periodic driving, while dynamical friction acts as dissipation. When the forcing frequency approaches the natural epicyclic frequency the system enters a resonance window where energy transfer from the soliton can overcome secular damping.
@@ III Eigenmode Decomposition
Our approach is based on Ref. [77], which we generalize to include the full set of angular states. Related eigenmode analyses have also been used to interpret soliton-core oscillations in FDM halos [42].The macroscopic ULDM wave function is equivalent to an effective single-particle wave function up to its normalization so we may construct eigenmodes using standard quantum-mechanical methods. In operator form the Schrödinger sector of the system is
@@ VII Conclusion
Several directions follow naturally. A first priority is to embed the soliton in a more realistic galactic environment by incorporating an outer halo, potentially using efficient wave-halo construction methods [76], and assessing the impact of ambient ULDM granule fluctuations, relaxation processes, and external tidal perturbations [9, 35, 73] on any dipole excitation of the soliton. Second, the resonance model should be generalized to eccentric or inclined orbits and multi-body configurations. Third, given that unequal-mass binaries or mergers could excite a dipole term it will be valuable to delineate the parameter space in which soliton-driven resonance works against a drag-driven inspiral. Fourth, one could generalize this work to models with a mixture of cold and ultralight dark matter [58] and multi-component ULDM models [62, 33, 5, 31]. Finally, we did not describe the original excitation of the dipole (and other) modes by the orbiting black hole(s) within perturbation theory – a full treatment of this would be both interesting and useful.
@@ B.2 Linearised quasi-circular expansion
$\ddot{\eta}=\frac{1}{r_{0}^{2}}$ $\Bigl[F_{\phi}(t)-2r_{0}\,\Omega_{0}\,\dot{\xi}$ (114)
$-\gamma_{\phi}\,r_{0}^{2}\,\bigl(\Omega_{0}+\dot{\eta}\bigr)-2\gamma_{\phi}\Omega_{0}r_{0}\xi\Bigr].$
@@ B.3 Forced damped harmonic oscillator
$\Delta$ $=\frac{16\bigl(2\Gamma^{2}-\kappa^{2}\bigr)^{2}}{9}-\frac{4\kappa^{4}}{3}$ (139)
$=\frac{4\bigl(\kappa^{2}-8\Gamma^{2}\bigr)^{2}-48\,\Gamma^{4}}{9}\geq 0,$
@@
$\mu_{0}$ $=0,$ (141)
$\qquad\mu_{\pm}$ $=\frac{2\kappa^{2}-4\Gamma^{2}\pm\sqrt{\kappa^{4}-16\kappa^{2}\Gamma^{2}+16\Gamma^{4}}}{3}.$